OpenStudy (anonymous):

what is the equation of the vertical line passing through the point (1, -1).

OpenStudy (jim766):

a vertical line is going to start with x= it passes through the point (1, -1) , the x being 1 so the equation would be x = 1

OpenStudy (jim766):

question?

OpenStudy (anonymous):

ah i love you thanks ! can you help with two more?

OpenStudy (jim766):

I can try

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4.

OpenStudy (jim766):

you can do this a couple different ways...do you want to use point slope form or y = mx + b?

OpenStudy (anonymous):

which is easier?

OpenStudy (jim766):

I like y = mx + b, but they are about the same amount of work. To write an equation of a line you need a slope and a y intercept. They gave you a slope in the problem, so all you need is the y intercept y = mx + b m is slope b is y intercept any of this familair?

OpenStudy (anonymous):

not really, im taking the pretest right now i havent been taught the lesson yet

OpenStudy (jim766):

in the point you gave do you know which is x and which is y?

OpenStudy (anonymous):

-3 x and -1 y

OpenStudy (jim766):

right, so just substitute them into y = mx + b -1 = 4(-3) + b can you solve this for b?

OpenStudy (jim766):

-1 = -12 +b +12 +12 11 = b right?

OpenStudy (jim766):

the m is 4, the b is 11 all you do now is subtitute those into y = mx + b y = 4x + 11

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

now look y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

those are my options so i dont have to solve it..

OpenStudy (jim766):

ok those are point slope form y - y1 = m(x - x1) y -(-1) = 4(x - 3) y + 1 = 4(x -3) its the first option

OpenStudy (anonymous):

thank you sooo much !

OpenStudy (anonymous):

but its y-1 not +1 right ?

OpenStudy (jim766):

the formula has a minus sign and the y is -1 y - - 1 = 4(x -3) minus a negative changes to addition y + 1 = 4(x -3)

OpenStudy (jim766):

make sense?

OpenStudy (anonymous):

i guess but its not one of the options

OpenStudy (anonymous):

its either y - 1 = 4(x - 3) or y + 1 = 4(x + 3)

OpenStudy (jim766):

sorry I lost the negative sign on the 3 on your point above it should be y + 1=4(x + 3) sorry for the confusion

OpenStudy (anonymous):

its ok thanks so much